OpenStudy (anonymous):

solve each system using substitution .. y=8-x and 7=2-y

OpenStudy (anonymous):

7= 2-8-x 7 = -6-x 13 = -x x = 13 y = 8-13 y = -4

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you are more than welcome